# Combining Crown Structures for Vulnerability Measures

### v1

@@ 2 Preliminaries
In this section, we briefly discuss terminology related to graphs and parameterized complexity that we use in the extended abstract (the complete terminology is given in LABEL:appendix::prelim). Additionally, we introduce the crown structures that are utilized throughout this paper.
Graph and Parameterized Terminology
We use a standard terminology for graphs $G=(V,E)$. The following is not necessarily common: For sets of vertex sets $\mathcal{V}\subseteq 2^{V}$ we abuse notation and use $V(\mathcal{V})=\bigcup_{Q\in\mathcal{V}}Q$ and $N(\mathcal{V}):=\left(\bigcup_{Q\in\mathcal{V}}N(U)\right)\setminus V(\mathcal{V})$. We define $\mathbb{CC}(G)\subseteq 2^{V}$ as the connected components of $G$ as vertex sets. Let $G=(V,E,w\colon V\to\mathds{N})$ be a vertex weighted graph. For $V^{\prime}\subseteq V$ and $G^{\prime}\subseteq G$ we define $w(V^{\prime})$ and $w(G^{\prime})$ as $\sum_{v\in V^{\prime}}w(v)$ and $\sum_{v\in V(G^{\prime})}w(v)$, respectively.
For parameterized complexity we use the standard terminology, which is also used, for example, in [downey2012parameterized, fomin2019kernelization]. Of particular interest in this work are kernelizations, which can be roughly described as formalized preprocessings. More formally, given an instance $(I,k)$ parameterized by $k$, a polynomial algorithm is called a kernelization if it maps any $(I,k)$ to an instance $(I^{\prime},k^{\prime})$ such that $(I^{\prime},k^{\prime})$ is a yes-instance if and only if $(I,k)$ is a yes-instance, $|I^{\prime}|\leq g(k)$, and $k^{\prime}\leq g^{\prime}(k)$ for computable functions $g,g^{\prime}$.
@@
We present a more general variant of crown decomposition that also captures commonly used crown compositions or expansions, which we explain in a moment. This novel decomposition can be easily derived from existing results of [DBLP:conf/esa/Casel0INZ21], but to the best of our knowledge has never been used in this form before. (See Figure 1 (left) for an illustration.)
Definition 2.1 (demanded balanced expansion and weighted crown decomposition).
For a graph $G=(A\cup B,E,w)$, a partition $A_{1}\cup A_{2}$ of $A$, an injective function $f\colon\mathbb{CC}(G[B])\to A$, demands $D=\{d_{a}\}_{a\in A}$ with $d_{a}\in\mathds{N}$ for each $a\in A$ and $y\in\mathds{N}$ the tuple $(A_{1},A_{2},y,f,D)$ is a demanded balanced expansion if
@@
If the demands are the same for each $a\in A$, e.g. $d_{a}=x$ for every $a\in A$ with $x\in\mathds{N}$, then we call $(A_{1},A_{2},y,f,x)$ a DBE. Furthermore, we call $(C,H,f)$ a ($x,y$) crown decomposition (($x,y$)-CD) if it is a $(H,\varnothing,y,f,x+y-1)$ DBE with $H=A$ and $C=f^{-1}(H)$.
@@
Left: Let $A=\{a_{1},a_{2},a_{3},a_{4}\}$ be ordered in a top down manner and $w(Q)=1$ for every $Q\in\mathbb{CC}(G[B])$. Then, ($(\{a_{1},a_{2}\},\{a_{3},a_{4}\},1,f,\{3,2,4,4\})$ is a DBE, where the assignment $f$ are depicted with corresponding colored bold edges.
Right: $\lambda$-balanced crown decomposition [DBLP:conf/esa/Casel0INZ21]
In Definition 2.1, the vertex set $A_{1}$ typically represents the structure targeted for reduction and can be seen as the ’head’ of a crown decomposition. The components within $\mathbb{CC}(G[B])$ assigned to $A_{1}$, denoted by $f^{-1}(A_{1})$, form the ’crown’. In the context of optimization problems such as the vertex cover problem, a successful reduction often involves the head being part of an optimal solution. Subsequently, upon its removal, the crown — separated by the head — no longer needs to be considered.
The classical crown decomposition corresponds to a $(1,1)$-CD, the $q$-expansion for $q\in\mathds{N}$ to a $(q,1)$-CD, the weighted crown decomposition for $x\in\mathds{N}$ to a $(x,x)$-CD, and for $x,y\in\mathds{N}$ a $(A_{1},A_{2},y,f,x)$ DBE is called a balanced expansion. The essential new structural property of our DBE are the varying demands for $A$.
Let $G=(A\cup B,E)$ be a graph. For $A^{\prime}\subseteq A$ we define $\mathcal{B}_{A^{\prime}}$ as the components $Q\in\mathbb{CC}(G[B])$ with $N(Q)\subseteq A^{\prime}$. The following theorem gives the runtime to find a DBE and the existence guarantee of $A_{1}$ depending on the size or weight in the graph.
Theorem 2.2 (demanded balanced expansion).
Let $G=\left(A\cup B,E,w\right)$ be a graph with no isolated components in $\mathbb{CC}(G[B])$, $y\geq\max_{Q\in\mathbb{CC}(G[B])}w(Q)$ and $D=\{d_{a}\}_{a\in A}$ demands with $d_{a}\in\mathds{N}$ for each $a\in A$. A demanded balanced expansion ($A_{1},A_{2},y,f,D$) can be computed in $\mathcal{O}\left(|V|\,|E|\right)$ time. Furthermore, if there is an $A^{\prime}\subseteq A$ with $w(A^{\prime})+w(V(\mathcal{B}_{A^{\prime}}))\geq\sum_{a\in A^{\prime}}d_{a}$, then $A_{1}\neq\varnothing$.
@@
Definition 2.3 ($\lambda$-balanced crown decomposition).
A $\lambda$-balanced crown decomposition of a graph $G=(V,E,w)$ is a tuple $((C,H,\mathfrak{R},f))$, where $\{H,C,R\}$ is a partition of $V$, the set $\mathfrak{R}$ is a partition of $R$, and $f\colon\mathbb{CC}(G[C])\rightarrow H$, such that:
@@
    $G[R^{\prime}]$ is connected and $\lambda<|R^{\prime}|\leq 3\lambda$ for each $R^{\prime}\in\mathfrak{R}$.
We use $\lambda$-BCD as an abbreviation for a $\lambda$-balanced crown decomposition. The authors [DBLP:conf/esa/Casel0INZ21] provide an algorithm that finds a $W$-BCD in polynomial time, as given in the following.
Theorem 2.4 (balanced crown decomposition theorem, [DBLP:conf/esa/Casel0INZ21] Theorem 2).
Let $G=(V,E,w)$ be a graph and $\lambda\in\mathds{N}$, such that each connected component in $G$ has weight larger than $\lambda$. A $\lambda$-balanced crown decomposition $(C,H,\mathfrak{R},f)$ of $G$ can be computed in $\mathcal{O}\left(r^{2}\,|V|\,|E|\right)$ time, where $r=|H|+|\mathfrak{R}|<|V|$ is at most the size of a maximum $(W+1)$-packing.
@@ 3 Improved Kernels for VI, wVI and wCOC
The vertex integrity of a graph models finding an optimal balance between removing vertices and keeping small connected parts of a graph. Expressed more formally: Given a graph $G=(V,E)$ and a parameter $p\in\mathds{N}$. The task is to find a vertex set $S\subseteq V$ such that $|S|+\max_{Q\in\mathbb{CC}(G-S)}|Q|$ is at most $p$. In the weighted vertex integrity problem (wVI), i.e. the vertices have weights, the aim is that $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)$ is at most $p$. We improve the vertex kernel of $p^{3}$ provided by Drange et al. [DBLP:journals/algorithmica/DrangeDH16] for both variants VI and wVI. To obtain such a kernel, they essentially established that vertices $v\in V$ with $w(v)+w(N(v))\geq p$ belong to a solution if we have a yes-instance in hand. A simple counting argument after removing those vertices provides a $p^{3}$ vertex kernel. We improve this outcome by employing crown decompositions, which offer valuable insights into the structural properties of vertex sets rather than focusing solely on individual vertices. Directly applying this method to problems like COC, as seen in prior works such as [DBLP:conf/esa/Casel0INZ21, DBLP:journals/tcs/ChenFSWY19, DBLP:conf/iwpec/KumarL16, DBLP:conf/tamc/XiaoK17], is challenging because we lack prior knowledge about the size of remaining components after solution removal. However, if we had at least a lower bound, we could theoretically safely reduce our instance accordingly. To establish such a bound, we engage in an interplay between packings and separators, where the balanced crown decomposition (BCD) proves instrumental. By identifying reducible structures in the input instance through embedding a DBE into BCD after determining a suitable bound, we prove the following theorems regarding VI and wVI.
Theorem 3.1.
@@
Theorem 3.3.
The weighted component order connectivity problem admits a vertex kernel of size $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. Furthermore, such a kernel can be computed in time $\mathcal{O}\left(r^{2}k|V|^{2}|E|\right)$, where $r$ is the size of a maximum $(W+1)$-packing.
Before we prove these theorems, we give some notations that we use in this section. We define them for the weighted case, as the unweighted case can be viewed in the same way with unit weights. We say that $S$ is a solution if it satisfies $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq p$ for wIV or $w(S)\leq k$ and $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq W$ for wCOC. We denote an instance of wIV by $(G,p)$ and an instance of wCOC by $(G,k,W)$. We say that $(G,p)$ or $(G,k,W)$ is a yes-instance if there is a solution, otherwise, we say that it is a no-instance. Let $\mathcal{S}\subseteq 2^{V(G)}$ be all solutions in $(G,p)$. We define $p_{\ell}:=\min_{S\in\mathcal{S}}\left(\max_{Q\in\mathbb{CC}(G-S)}w(Q)\right)$, which is a lower bound for any solution $S\in\mathcal{S}$ on the size of the connected components after its removal. For all instances $(G,p)$ of wVI, we assume that $w(v)<p$ for every $v\in V$ and that $G$ contains a connected component of weight more than $p$. Otherwise, the instance is trivial.
@@ 3.1 Vertex Integrity
The rest of this section is dedicated to the proof of Theorem 3.1. First, we discuss the reducible structure we are looking for and then show how we can find it.
@@
Let $(G,p)$ be an instance of VI and let $(C,H,f)$ be a $(p,p_{\ell})$-CD in $G$. Then, $(G,p)$ is a yes-instance if and only if $(G-\{H\cup C\},p-|H|)$ is a yes-instance.
In order to use Lemma 3.4, we must first determine a suitable lower bound for $p_{\ell}$. Additionally, we need to ensure the existence of a corresponding weighted crown decomposition when the input graph size exceeds $3p^{2}$, and that we can find it efficiently. This is where the BCD comes into play. For $\lambda\in\mathds{N}$ a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ in $G$ can only be computed within the components of $G$ that have a size (or weight) greater than $\lambda$. For a better readability, we will consistently assume that when computing a $\lambda$-BCD, we disregard small components, meaning that if $(C,H,\mathfrak{R},f)$ is a $\lambda$-BCD for $G$, then $C\cup H\cup V(\mathfrak{R})$ are only the vertices that are contained in a component of size (or weight) more than $\lambda$ of $G$.
We specify the following result directly in the weighted version to also use it later.

### v2

@@ 2 Preliminaries
In the upcoming paragraphs, we discuss the standard terminology related to graphs and parameterized complexity. Additionally, we introduce some crown structures that are utilized throughout this paper.
Graph Terminology
Let $G=(V,E)$ be a graph. We use $V(G)$ and $E(G)$ to denote $V$ and $E$, respectively. We define the size of a subgraph $G^{\prime}\subseteq G$ as the number of its vertices, where we denote the size of $G$ by $n$. For $v\in V(G)$ we denote its neighborhood by $N(v):=\{u\in V(G):uv\in E(G)\}$. Let $V^{\prime}\subseteq V$ be a vertex subset. We define $G[V^{\prime}]$ as the induced subgraph of $V^{\prime}$, $G-V^{\prime}:=G[V\setminus V^{\prime}]$ and $N(V^{\prime}):=\left(\bigcup_{v\in V^{\prime}}N(v)\right)\setminus V^{\prime}$. For a vertex $v$ we define $G+v$ as the graph $G$ with an additional isolated vertex $v$. Similarly, for a graph $H$ we define $G+H$ as the graph $G$ with an additional isolated graph $H$. For sets of vertex sets $\mathcal{V}\subseteq 2^{V}$ we abuse notation and use $V(\mathcal{V})=\bigcup_{Q\in\mathcal{V}}Q$ and $N(\mathcal{V}):=N(V(\mathcal{V}))=\left(\bigcup_{Q\in\mathcal{V}}N(Q)\right)\setminus V(\mathcal{V})$. We define $\mathbb{CC}(G)\subset 2^{V}$ as the connected components of $G$ as vertex sets. For vertex subsets $H,C\subset V$, we define a function $f\colon\mathbb{CC}(C)\to H$. For any subset $H^{\prime}\subseteq H$, we frequently use the inverse $f^{-1}(H^{\prime})$, which returns sets of vertices. Typically, however, we are interested in the union of these vertex sets, denoted $V(f^{-1}(H^{\prime}))$. For readability, we may omit the vertex operator when it is clear from the context that we refer to the union. We use the same terminology for vertex weighted graphs, denoted by $G=(V,E,w)$ where $w$ is a weight function $w\colon V\to\mathds{N}$. Let $G=(V,E,w)$ be a vertex weighted graph. For $V^{\prime}\subseteq V$, $G^{\prime}\subseteq G$ (meaning that $G^{\prime}$ is a subgraph of $G$) and $\mathcal{V}\subset 2^{V}$ we define $w(V^{\prime})$, $w(G^{\prime})$ and $w(G^{\prime})$ as $\sum_{v\in V^{\prime}}w(v)$, $\sum_{v\in V(G^{\prime})}w(v)$ and $\sum_{v\in V(\mathcal{V})}w(v)$, respectively.
Parameterized Terminology
We use the standard terminology for parameterized complexity, which is also used, for example, in [downey2012parameterized, fomin2019kernelization]. A parameterized problem is a decision problem with respect to certain instance parameters. Instances of a parameterized problem are usually given as tuple $(I,k)$, where $k$ is the parameter.
If there exists an algorithm that decides each instance $(I,k)$ of a parameterized problem $\Pi$ in time $f(k)\cdot|I|^{c}$, where $f$ is a computable function and $c$ is a constant, then $\Pi$ is called fixed-parameter tractable. We say $(I,k)$ is a yes-instance if the answer to the decision problem is positive, otherwise we say $(I,k)$ is a no-instance.
Of particular interest in this work are kernelizations, which can be roughly described as formalized preprocessing. More formally, a polynomial algorithm is called a kernelization for a parameterized problem $\Pi$ if it maps any instance $(I,k)$ of $\Pi$ to an instance $(I^{\prime},k^{\prime})$ of $\Pi$ such that $(I^{\prime},k^{\prime})$ is a yes-instance if and only if $(I,k)$ is a yes-instance, $|I^{\prime}|\leq g(k)$, and $k^{\prime}\leq g^{\prime}(k)$ for computable functions $g,g^{\prime}$.
@@
We present a more general variant of crown decomposition that also captures commonly used crown decompositions or expansions (strictly speaking the relevant parts of it), which we explain in a moment. This novel decomposition can be easily derived from existing results of [DBLP:conf/esa/Casel0INZ21], but to the best of our knowledge has never been used in this form before (see Fig. 1 (left) for an illustration).
Definition 1 (Demanded balanced expansion and weighted crown decomposition).
For a graph $G=(A\cup B,E,w)$, a partition $A_{1}\cup A_{2}$ of $A$, a function $f\colon\mathbb{CC}(G[B])\to A$, demands $D=\{d_{a}\}_{a\in A}$ with $d_{a}\in\mathds{N}$ for each $a\in A$ and $y\in\mathds{N}$ the tuple $(A_{1},A_{2},y,f,D)$ is a demanded balanced expansion (DBE) if
@@
To simplify the notation we introduce two further special cases of a DBE:
  • If the demands are the same for each $a\in A$, e.g. $d_{a}=x$ for every $a\in A$ with $x\in\mathds{N}$, then we write only the value $x$ instead of a vector $D=\{d_{a}\}_{a\in A}$ in a DBE-tuple, i.e. $(A_{1},A_{2},y,f,D)=(A_{1},A_{2},y,f,x)$.
  • For $q\in\mathds{N}$ we call $(C,H,f)$ a ($q,y$) crown decomposition (($q,y$)-CD) if it is a $(H,\varnothing,y,f,q+y-1)$ DBE with $H=A$ and $C=B=f^{-1}(H)$. (This term simplifies the reference to the reducible structure of crown $C$ and head $H$.)
@@
Left: Let $A=\{a_{1},a_{2},a_{3},a_{4}\}$ be ordered in a top down manner, $w(Q)=1$ for every $Q\in\mathbb{CC}(G[B])$ and $w(a)=1$ for every $a\in A$. Then, ($(\{a_{1},a_{2}\},\{a_{3},a_{4}\},1,f,\{3,1,3,4\})$ is a DBE, where the assignment $f$ are depicted with corresponding colored bold edges.
Right: A $\lambda$-balanced crown decomposition, where the assignment $f$ are depicted with corresponding colored bold edges. The two dashed lines illustrate that $w(h)+w(f^{-1}(h))>\lambda$ for every $h\in H$ while $w(Q)\leq\lambda$ for every $Q\in\mathbb{CC}(G[C])$.
Considering a graph instance $G=(V,E)$, e.g. from VI or COC, Definition 1 usually describes only a subgraph where the reducible structure is sought, i.e. $A,B$ are vertex subsets of $V$. It is crucial that $A$ already separates $B$ from $V\setminus B$. Then, the vertex set $A_{1}\subseteq A$ (cf. Definition 1) typically represents the structure targeted for reduction and can be seen as the “head” of a crown decomposition. The components within $\mathbb{CC}(G[B])$ assigned to $A_{1}$ (the separated part through $A_{1}$), denoted by $f^{-1}(A_{1})$, form the “crown”. In the context of optimization problems such as the vertex cover problem, a successful reduction often involves the head being part of an optimal solution. Subsequently, upon its removal, the crown — separated by the head — no longer needs to be considered.
Ideally, we wish to have $A_{1}=A$, but even when this is not the case the balanced part of a DBE ensures that the elements mapped to $A_{2}$ are bounded, which finally allows us to bound the number of vertices (or the sum of the vertex weights) of $A_{2}\cup V(f^{-1}(A_{2}))$.
Consider a crown decomposition, which is a partition of the vertex set into body, head and crown, where the head separates the body from the crown. The head and crown of the classical crown decomposition corresponds to a $(1,1)$-CD, the $q$-expansion for $q\in\mathds{N}$ to a $(q,1)$-CD and the weighted crown decomposition for $q\in\mathds{N}$ to a $(q,x)$-CD. The last known structure captured by Definition 1 is the balanced expansion which corresponds for $x,y\in\mathds{N}$ to a $(A_{1},A_{2},y,f,x)$ DBE. The essential new structural property of a DBE compared to a balanced expansion are the varying demands for $A$.
Let $G=(A\cup B,E)$ be a graph. For $A^{\prime}\subseteq A$ we define $\mathcal{B}_{A^{\prime}}$ as the components $Q\in\mathbb{CC}(G[B])$ with $N(Q)\subseteq A^{\prime}$. The following theorem, easily derived as modification of the results of [DBLP:conf/esa/Casel0INZ21], gives the runtime to find a DBE and the existence guarantee of $A_{1}$ depending on the size or weight in the graph. For the sake of completeness, we give the proof in Appendix LABEL:app:a.
Theorem 1 (Demanded balanced expansion).
Let $G=\left(A\cup B,E,w\right)$ be a graph with no isolated components in $\mathbb{CC}(G[B])$, i.e. every component of $\mathbb{CC}(G[B])$ contains at least one neighbor of $A$. Let $y\geq\max_{Q\in\mathbb{CC}(G[B])}w(Q)$ and $D=\{d_{a}\}_{a\in A}$ demands with $d_{a}\in\mathds{N}$ for each $a\in A$. A demanded balanced expansion ($A_{1},A_{2},y,f,D$) can be computed in $\mathcal{O}\left(|V|\,|E|\right)$ time. Furthermore, if there is an $A^{\prime}\subseteq A$ with $w(A^{\prime})+w(V(\mathcal{B}_{A^{\prime}}))\geq\sum_{a\in A^{\prime}}d_{a}$, then $A_{1}\neq\varnothing$.
@@
Definition 2 ($\lambda$-balanced crown decomposition).
A $\lambda$-balanced crown decomposition of a graph $G=(V,E,w)$ is a tuple $(C,H,\mathfrak{R},f)$, where $\{H,C,R\}$ is a partition of $V$, the set $\mathfrak{R}$ is a partition of $R$, and $f\colon\mathbb{CC}(G[C])\rightarrow H$, such that:
@@
    $G[R^{\prime}]$ is connected and $\lambda<w(R^{\prime})\leq 3\lambda$ for each $R^{\prime}\in\mathfrak{R}$.
We use $\lambda$-BCD as an abbreviation for a $\lambda$-balanced crown decomposition. Looking at the original definition of a $\lambda$-BCD in [DBLP:conf/esa/Casel0INZ21] (Definition 6), we shift the $\lambda$ value by one, which allows us to change several inequalities between strict and non-strict for a clearer representation, while still keeping the definition the same. Furthermore, $\lambda$ must be at least two in the original definition, but to simplify the understanding of the application to the problems considered in this paper, it makes more sense that a $\lambda$-BCD can also exist with $\lambda=1$.
The authors in [DBLP:conf/esa/Casel0INZ21] provide an algorithm that finds a $\lambda$-BCD in polynomial time, as given in the following.
Theorem 2 (Balanced crown decomposition theorem, [DBLP:conf/esa/Casel0INZ21] Theorem 7).
Let $G=(V,E,w)$ be a graph and $\lambda\in\mathds{N}$, such that each connected component in $G$ has weight larger than $\lambda$. A $\lambda$-balanced crown decomposition $(C,H,\mathfrak{R},f)$ of $G$ can be computed in $\mathcal{O}\left(r^{2}\,|V|\,|E|\right)$ time, where $r=|H|+|\mathfrak{R}|<|V|$ is at most the size of a maximum $(\lambda+1)$-packing.
@@ 3 Improved Kernels for VI, wVI and wCOC
The vertex integrity of a graph models finding an optimal balance between removing vertices and keeping small connected parts of a graph. As a reminder: In the formal definition, a graph $G$ and a parameter $p$ are given. The task is to find a vertex set $S\subseteq V$ such that $|S|+\max_{Q\in\mathbb{CC}(G-S)}|Q|$ is at most $p$. In the weighted vertex integrity problem (wVI), i.e. the vertices have weights, the aim is that $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)$ is at most $p$. We improve the vertex kernel of $p^{3}$ provided by Drange et al. [DBLP:journals/algorithmica/DrangeDH16] for both variants VI and wVI. To obtain such a kernel, they essentially established that vertices $v\in V$ with $w(v)+w(N(v))>p$ belong to a solution if we have a yes-instance in hand. A simple counting argument after removing those vertices provides a $p^{3}$ vertex kernel. We improve this by employing crown decompositions, which offer valuable insights into the structural properties of vertex sets rather than focusing solely on individual vertices. Directly applying this method to problems like COC, as seen in prior works such as [DBLP:conf/esa/Casel0INZ21, DBLP:journals/tcs/ChenFSWY19, DBLP:conf/iwpec/KumarL16, DBLP:conf/tamc/XiaoK17], is challenging because we lack prior knowledge about the size of remaining components after solution removal. However, if we had at least a lower bound, we could theoretically safely reduce our instance accordingly. To establish such a bound, we engage in an interplay between packings and separators, where the balanced crown decomposition (BCD) proves instrumental. By identifying reducible structures in the input instance through embedding a DBE into BCD after determining a suitable bound, we prove the following theorems regarding VI and wVI.
Theorem 3.
@@
Theorem 5.
The weighted component order connectivity problem admits a vertex kernel of size $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. Furthermore, such a kernel can be computed in time $\mathcal{O}\left(r^{2}k|V||E|\right)$, where $r$ is the size of a maximum $(W+1)$-packing.
Before we prove these theorems, we give some notations that we use in this section. We define them for the weighted case, as the unweighted case can be viewed in the same way with unit weights. We say that $S$ is a solution if it satisfies $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq p$ for wVI or $w(S)\leq k$ and $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq W$ for wCOC. We denote an instance of wVI by $(G,p)$ and an instance of wCOC by $(G,k,W)$. We say that $(G,p)$ or $(G,k,W)$ is a yes-instance if there is a solution, otherwise, we say that it is a no-instance. Let $\mathcal{S}\subseteq 2^{V(G)}$ be all solutions for $(G,p)$. We define $p_{\ell}:=\min_{S\in\mathcal{S}}\left(\max_{Q\in\mathbb{CC}(G-S)}w(Q)\right)$ which is the optimum lower bound on the size of the connected components after the removal of any solution; where for no-instances we set $p_{\ell}=p$. For all instances $(G,p)$ of wVI, we assume $w(v)<p$ for every $v\in V$ and that $G$ contains a connected component of weight more than $p$. Lastly, for a better readability, we will consistently assume that when computing a $\lambda$-BCD, we disregard small components, meaning that if $(C,H,\mathfrak{R},f)$ is a $\lambda$-BCD for $G$, then $C\cup H\cup V(\mathfrak{R})$ are only the vertices that are contained in some component of size (or weight) more than $\lambda$ of $G$.
@@ 3.1 Vertex Integrity
The rest of this section is dedicated to the proof of Theorem 3. We first give the types of reductions we want to use, always with respect to some $c\leq p_{\ell}$, and then discuss how to efficiently compute an estimate of $c$ to apply these reduction rules.
Lemma 6.
Let $(G,p)$ be an instance of VI where not all connected components are of size at most $p$ and let $(C,H,f)$ be a $(p,c)$-CD in $G$ for some $c\leq p_{\ell}$, such that $N(C)\subseteq H$. Then, $(G,p)$ is a yes-instance if and only if $(G-(H\cup C)+K_{c},p-|H|)$ is a yes-instance.
Proof 3.7.
Let $G^{\prime}=G-\{H\cup C\}+K_{c}$. First assume $(G,p)$ is a yes-instance, and let $S$ be an according solution with $s=\max_{Q\in\mathbb{CC}(G-S)}|Q|$. Note that since not all components of $G$ have size at most $p$, it follows that $S\neq\emptyset$ and $s<p$. We claim that $S^{\prime}=S\setminus(H\cup C)$ is a solution for $(G^{\prime},p-|H|)$. It suffices to show that $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}|Q|\leq s$ and $|S^{\prime}|\leq|S|-|H|$. More precisely, we can then conclude that $S^{\prime}$ has size at most $|S^{\prime}|\leq(p-|H|)-s$ and creates components of size at most $s$ when deleted from $G^{\prime}$, which makes it a solution for $(G^{\prime},p-|H|)$, meaning that $(G^{\prime},p-|H|)$ is a yes-instance.
For the claim $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}|Q|\leq s$, first note that $c\leq p_{\ell}$ by the choice of $c$, and that by the definition of $p_{\ell}$ we have $p_{\ell}\leq s$. Thus the $K_{c}$ component of $G^{\prime}$ has size at most $s$. For the other components in $G^{\prime}-S^{\prime}$, suppose that $G-(H\cup C)-S^{\prime}$ contains a connected component $Q^{\prime}$ of size larger than $s$. Note that $V(Q^{\prime})\subset V\setminus(H\cup C)$. Since we do not remove vertices from $S$ in $V\setminus(C\cup H)$ to build $S^{\prime}$, we have that $Q^{\prime}$ is also a component of $G-S$, contradicting $s=\max_{Q\in\mathbb{CC}(G-S)}|Q|$.
For the claim $|S^{\prime}|\leq|S|-|H|$ observe that $\mathcal{V}_{H}=\{\{h\}\cup V(f^{-1}(h))\}_{h\in H}$ forms a $p$-packing of size $|H|$ in $G[C\cup H]$. Since $p>s$ and the elements in $\mathcal{V}_{H}$ are pairwise disjoint, every element of $\mathcal{V}_{H}$ contains at least one vertex of $S$. Thus, we obtain $|S\cap(C\cup H)|\geq|H|$ and therefore $|S^{\prime}|\leq|S|-|H|$.
For the other direction, assume $(G^{\prime},p-|H|)$ is a yes-instance and let $S^{\prime}$ be a solution for $(G^{\prime},p-|H|)$, and let $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}|Q|=s^{\prime}$. We claim that $S=(S^{\prime}\cup H)\cap V$ is a solution for $(G,p)$.
Components of $G-S$ are either the components of $G[C\cup H]-H$, since $H$ separates $C$ from the rest of $G$, or components of $G^{\prime}-S^{\prime}$. By definition of the crown decomposition, all components of $G[C\cup H]-H$ are of size at most $c$, while all components of $G^{\prime}-S^{\prime}$ are of size at most $s^{\prime}$ by the definition of $s^{\prime}$. If $c\leq s^{\prime}$ it follows that all components of $G-S$ have size at most $s^{\prime}\leq p-|H|-|S^{\prime}|\leq p-|S|$ (since $S^{\prime}$ is a solution for $(G^{\prime},p-|H|)$ creating components of size $s^{\prime}$, meaning that $s^{\prime}+|S^{\prime}|\leq p-|H|$, and $|S|\leq|S^{\prime}|+|H|$), and we are done.
Assume that $s^{\prime}<c$. This means that the $K_{c}$ component in $G^{\prime}$ is shrunk to have size at most $s^{\prime}$, meaning that $S^{\prime}$ contains at least $c-s^{\prime}$ vertices from $K_{c}$. This gives $|S|\leq|S^{\prime}|+|H|-c+s^{\prime}$. If $G-S$ now contains a component larger than $s^{\prime}$, this component originates from the crown and has a size of at most $c$. This ensures that $S$ is a solution for $(G,p)$ because
$|S|+c$ $\leq|S^{\prime}|+|H|-c+s^{\prime}+c$
$=|S^{\prime}|+s^{\prime}+|H|$
$\leq p-|H|+|H|$
$\leq p.$
We will eventually use BCDs to find a good estimate for $c$ and to find a CD that can be removed with Lemma 6. Since we only search for $\lambda$-BCDs in components of size larger than $\lambda$, we still need another reduction rule for small components. Clearly, if an instance contains more than $p$ components of size at most $c$, deleting a smallest such component is a safe rule. With the same trick to add a $K_{c}$ like in Lemma 6 we can go further and delete all other small components.
@@
Let $(G,p)$ be an instance of VI, $c<p_{\ell}$, and let $U$ be a connected component of $G$ of size at most $c$. Then $(G,p)$ is a yes-instance if and only if $(G-U+K_{c},p)$ is a yes-instance.
Proof 3.9.
If $(G,p)$ is a yes-instance, let $S$ be a corresponding solution, meaning that for $s=\max_{Q\in\mathbb{CC}(G-S)}|Q|$, we have $|S|+s\leq p$. By the definition of $p_{\ell}$, we know that $s\geq p_{\ell}$. Each component of $G-U+K_{c}-S$ is a component of $G-S$, and thus has size at most $s$, and the $K_{c}$, which has size $c\leq p_{\ell}\leq s$. Therefore, $S$ is also a solution for $(G-U+K_{c},p)$.
For the other direction, assume $(G-U+K_{c},p)$ is a yes-instance and let $S$ be such that $s=\max_{Q\in\mathbb{CC}(G-U+K_{c}-S)}|Q|$ satisfies $|S|+s\leq p$. We claim that $S^{\prime}=S\cap V(G)$ is a solution for $(G,p)$. Since $U$ is a connected component of $G$, all components in $G-S^{\prime}$, except for $U$, are components of $G-U+K_{c}-S$, so they have size at most $p-|S|\leq p-|S^{\prime}|$. Thus if $|U|\leq s$ then it immediately follows that all components of $G-S^{\prime}$ have size at most $p-|S^{\prime}|$ meaning that $S^{\prime}$ is a solution for $(G,p)$.
If $|U|>s$ then $U$ is a largest component in $G-S^{\prime}$. Since $|U|\leq c$, it follows that $c>s$ which means that the $K_{c}$ component in $G-U+K_{c}$ must be shrunk by $S$. More precisely, $S$ contains at least $c-s$ of its vertices. This means $|S^{\prime}|\leq|S|-c+s$ which gives $|U|+|S^{\prime}|\leq c+|S|-c+s=|S|+s\leq p$. With $U$ being a largest component in $G-S^{\prime}$ it again follows that $S^{\prime}$ is a solution for $(G,p)$.
Note that exhaustive application of Lemma 3.8 creates an equivalent instance with only one component of size at most $c$ (isomorphic to $K_{c}$).
In order to use our two reduction rules, we must first determine a suitable lower bound for $p_{\ell}$. Additionally, we need to ensure the existence of a corresponding weighted crown decomposition when the input graph size exceeds a certain size, and that we can find it efficiently. This is where the BCD comes into play. For $\lambda\in\mathds{N}$ a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ in $G$ can only be computed within the components of $G$ that have a size (or weight) greater than $\lambda$. As a reminder, when we compute a $\lambda$-BCD, we disregard small components, meaning that if $(C,H,\mathfrak{R},f)$ is a $\lambda$-BCD for $G$, then $C\cup H\cup V(\mathfrak{R})$ are only the vertices that are contained in some component of size more than $\lambda$ of $G$.
We specify the following lemma directly in the weighted version to also use it later. The lemma provides a lower bound for $p_{\ell}$ from a $\lambda$-BCD $(C,H,\mathfrak{R},f)$, where we essentially use that a $\lambda$-BCD is also a $(\lambda+1)$-packing of size $|H|+|\mathfrak{R}|$.